OpenStudy (anonymous):

What is the solution to the equation √2x+8− 6 = 4

OpenStudy (anonymous):

added 6 to both sides to make 4 a 10

OpenStudy (raden):

yeah, that's right now you have √2x+8 = 10 then square both sides, what do you get ?

OpenStudy (anonymous):

2x + 8 = sqrt of 10?

OpenStudy (anonymous):

im stuck

OpenStudy (raden):

a bit mistake in right side, 10^2 = 100

OpenStudy (raden):

so, that should be 2x + 8 = 100

OpenStudy (anonymous):

so its x=46

OpenStudy (raden):

yes, i got like that too :)

OpenStudy (anonymous):

thanks

OpenStudy (raden):

you're welcome